If 2.83 m3 of atmospheric air (101.3 kPa) at 0oC are compressed to a volume of 0.0283 m3 at a temperature of 76oC, what will be the pressure of the air in kPa.


1
Expert's answer
2021-12-09T09:15:16-0500

"\frac{p_1V_1}{T_1}=\frac{p_2V_2}{T_2},\implies"

"p_2=p_1\frac{V_1T_2}{V_2T_1}=12.95~MPa."